# Watermarking Algorithm for Large Language Models

### v1

@@ 3.2 Watermarked Text Detection
The detection of SimMark follows a similar methodology to KGW by employing a one-proportion $z$-test for hypothesis testing. However, akin to Hou et al. (2023, 2024), the detection operates at the sentence level rather than the token level. To perform detection, we first divide the input text into sentences [^2] Using sent_tokenize method of NLTK (Bird et al., 2009). and use the same Instructor-Large embedding model with identical instruction to compute the embeddings for each sentence. If PCA was used during the watermark generation process, it must also be applied here to ensure consistency.
Algorithm 2 Detection Pseudo-Code for SimMark
0:  input text, embedding model, predefined interval $[a,b]$, decay factor $K$, threshold $\beta$, optional PCA model
1:  Split the input text into sentences excluding the first sentence (i.e, the prompt): $M=\{X_{2},\dots,X_{N_{\mathrm{total}}}\}$.
@@
8:  $N\leftarrow|M|$
9:  Soft count of valid sentences: $N_{\mathrm{valid\_soft}}\leftarrow\sum_{i=2}^{N+1}c_{i}$.
@@
10:  Estimate $p_{0}$ as the area under the human-written text embeddings similarity distribution curve within $[a,b]$.
11:  Compute $z_{\mathrm{soft}}$ (i.e., the soft-$z$-score) using Eq. (2).
12:  if $z_{\mathrm{soft}}>\beta$ then
@@
13:     Reject $H_{0}$, i.e., the input text is likely generated by an LLM watermarked by SimMark.  [Uncaptioned image]
14:  else
15:     Accept $H_{0}$, i.e., the input text is likely written by a human.   [Uncaptioned image]
16:  end if
@@
Next, we compute the similarity (or distance) between consecutive sentences and count the number of valid sentences, denoted as $N_{\mathrm{valid\_soft}}$. A sentence $X_{i+1}$ is deemed valid if $s_{i+1}$ the similarity (or distance) between $X_{i+1}$ and $X_{i}$ lies within the predefined user interval $[a,b]$. However, paraphrasing attacks may alter embeddings significantly, causing the similarity (or distance) to deviate from the desired interval. To mitigate this, we adopt a soft counting approach, where a sentence is considered partially valid if its similarity is near the interval boundaries. Specifically, the soft count of $X_{i+1}$, denoted as $c_{i+1}$, is defined as follows:
$c_{i+1}=c(s_{i+1})=\begin{cases}1&\text{if }s_{i+1}\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\in[a,b]},\\ e^{-K\min\{|a-s_{i+1}|,|b-s_{i+1}|\}}&{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\text{otherwise.}}\end{cases}\end{aligned}$
@@
Here, $s_{i+1}=sim(e_{i},e_{i+1})$ represents the similarity (or distance) between consecutive sentence embeddings, and $K>0$ is a decay factor controlling the smoothness of the soft counting mechanism. A higher $K$ makes the function behave closer to a step function, while a lower $K$ allows for smoother transitions, tolerating minor deviations outside the interval $[a,b]$. The total number of valid sentences is then computed as $N_{\mathrm{valid\_soft}}=\sum c_{i}$.
Refer to Appendix C for experimental results demonstrating how this soft counting mechanism improves robustness against paraphrasing, with only a minimal impact on performance in non-paraphrased scenarios. Allowing for some degree of error through soft counting enables better detection under adversarial conditions while maintaining high accuracy in non-paraphrased cases.
@@
During our initial experiments, we observed that contrary to cosine similarity, Euclidean distance is very sensitive to paraphrasing and even a subtle change in the sentences would result in a huge difference in the distances of embeddings. We hypothesize that Euclidean distance is sensitive to noise in high-dimensional spaces such as the semantic space of the embedder. To mitigate this, we propose using PCA: We first train a PCA model on a dataset of human-written texts, to find the principal components of the sentence embeddings, i.e., the components that contribute the most to the semantical representation of the sentences. Then, we apply PCA to the embeddings to reduce their dimension. More details on the dimensionality reduction are provided in Section 4 and Appendix D.
With this approach reverse-engineering the watermark algorithm becomes even harder as the third party who tries to reverse-engineer the algorithm, not only requires to know the embedder model and the instruction used for it, but it also needs to know the specifics of PCA setting (such as the number of principal components) and needs access to the dataset for training the PCA; otherwise, the similarity distribution of sentence embeddings would be different, making reverse engineering challenging. We found that this approach improved the robustness against paraphrasing when Euclidean distance is used. However, it would decrease the performance when cosine similarity is used.
Returning to the detection, the null hypothesis $H_{0}$ is defined as follows:
@@
$H_{0}$: The sentences are written by humans, i.e., the text sequence is written (or generated) without knowledge of the valid interval in the similarity distribution of sentence embeddings.
@@
We calculate the $z$-statistic for the one-proportion $z$-test using the sample proportion $p=\frac{N_{\mathrm{valid\_soft}}}{N}$, where $N$ is the total number of samples (sentences). Since $N_{\mathrm{valid\_soft}}$ is a soft count of valid sentences, we refer to it as soft-$z$-score, which is given by $z_{\mathrm{soft}}=\frac{p-p_{0}}{\sqrt{\frac{p_{0}(1-p_{0})}{N}}}$. Here, $p_{0}$ is the population proportion and represents the ratio of valid sentences to all sentences in human-written text (i.e., in a text with no watermark), which is estimated as the area under the similarity distribution curve of consecutive human-written sentences within the interval $[a,b]$ (Like the one in Figure 6). Alternatively, the soft-$z$-score can be expressed as:
@@
In this form, the value of $z_{\mathrm{soft}}$ can be interpreted as a normalized deviation of the number of valid sentences $N_{\mathrm{valid\_soft}}$ from its expectation $p_{0}N$. As highlighted in Algorithm 2, the null hypothesis $H_{0}$ is rejected if $z_{soft}>\beta$, where $\beta$ is a threshold determined empirically by running the detection algorithm on human-written text. The threshold $\beta$ is selected to maintain a desired FP rate (i.e., minimizing the misclassification of human-written text as LLM-generated). Details on the computation of $\beta$ are provided in the Appendix F.
4 Experiments
@@
[…]
No watermark 11.89 11.43 3.10
SemStamp 12.89 11.51 3.17
$k$-SemStamp 11.82 11.48 3.11
Cosine-SimMark (Ours) 12.69 11.50 3.39
Euclidean-SimMark [^7] PCA is applied. (Ours) 9.67 11.50 3.28
Table 2: Comparison of the quality of text generated using different watermarking algorithms on the BookSum dataset. For PPL metric, lower values indicate better quality, while higher values are better for Ent-3 and Sem-Ent. SimMark watermarking achieves quality comparable to no watermarking, demonstrating minimal impact on text quality.
In this section, we evaluate the performance of our watermarking algorithm, SimMark, across different datasets. Performance is assessed using the area under the receiver operating characteristic curve (ROC-AUC) and true positive rate (TP) at fixed false positive rates (FP) of 1% and 5%, (TP@1%FP and TP@5%FP). Higher values indicate better performance across all metrics.
For dimensionality reduction, we trained a PCA model on 8000 samples from the RealNews subset of the C4 dataset (Raffel et al., 2020), reducing embedding dimensions from 768 to 16. After testing various principal component counts (ranging from 512 to 16), we found 16 to yield the best results. PCA improved robustness against paraphrasing attacks when Euclidean distance was used (except on the BookSum dataset) but consistently degraded performance when cosine similarity was employed across all datasets. During our experiments, we applied PCA wherever it enhanced performance. More details and the results of these experiments are summarized in Table 4 in Appendix D.
Across all experiments, the decay factor was set to $K=250$, as this value provided an optimal trade-off between performance under both non-paraphrased and paraphrased conditions (See Appendix C). The threshold $\beta$ was determined programmatically during the detection (See Appendix F) to achieve the specified FP rates (1% or 5%). The predefined intervals $[0.68,0.76]$ for cosine similarity and $[0.28,0.36]$ for Euclidean distance with PCA and $[0.4,0.55]$ for Euclidean distance without PCA were found to be near-optimal. While the intervals could have been further optimized for each dataset individually, we chose not to do so to demonstrate the general performance of our method. For additional details on the experimental environment refer to Appendix H.
@@ 4.1 Models and Datasets
To ensure comparability with prior works by Hou et al. (2023, 2024), we also employed the same fine-tuned version of OPT-1.3B (Zhang et al., 2022), that they used as our LLM [^3] Used AbeHou/opt-1.3b-semstamp model.. For semantic embedding, we utilized the Instructor-Large model [^4] Used hkunlp/instructor-large embedding model. (Su et al., 2023), as mentioned earlier. Following Kirchenbauer et al. (2023), sampling from the LLM was performed with a temperature of 0.7 and a repetition penalty of 1.05, while the minimum and the maximum number of generated tokens were set to 195 and 205, respectively. The maximum number of iterations for rejection sampling is set to 100, following Hou et al. (2023, 2024).
In our experiments, we used the RealNews subset of the C4, the BookSum (Kryscinski et al., 2022), and Reddit-TIFU (Kim et al., 2019) datasets, as in Hou et al. (2023). 1000 samples from each dataset were chosen to analyze the detection results and the text quality. Each sample was segmented into sentences, with the first sentence serving as the prompt to the LLM.
@@
  • Perplexity (PPL): Measures how surprising the text is to a larger LLM (OPT-2.7B (Zhang et al., 2022)). Lower values indicate higher quality.
@@
  • Tri-gram Entropy (Ent-3) (Zhang et al., 2018): To assess textual diversity via the entropy of the tri-grams distribution.
  • Semantic Entropy (Sem-Ent) (Han et al., 2022): Used to measure semantic informativeness and diversity of the text.
4.2 Paraphrasing Attack
@@
To evaluate the robustness of SimMark, we tested it against paraphrasing attacks using three models: Pegasus paraphraser [^5] Used tuner007/pegasus_paraphrase paraphrasing model. (Zhang et al., 2020), Parrot paraphraser (Damodaran, 2021), and GPT-3.5-Turbo (OpenAI, 2022). Kirchenbauer et al. (2024) observed that prompting models to paraphrase entire texts often results in summarized outputs, with the summarization ratio worsening for longer inputs. To prevent such summarization, we adopted a sentence-by-sentence paraphrasing scheme, which also ensures our results are comparable to prior works such as Hou et al. (2023, 2024). The quality of paraphrases was assessed using BertScore [^6] Used microsoft/deberta-xlarge-mnli (He et al., 2021) model. (Zhang* et al., 2020), with all settings consistent with prior works of Hou et al. (2023, 2024). The bottom part of Figure 2 demonstrates the paraphrase attack and detection phase in more detail. Refer to Appendix I to find the prompts used with GPT-3.5-Turbo.
@@
We also included the results for the bigram paraphrase attacks introduced by Hou et al. (2023), with identical settings (25 rephrases for each sentence, etc.). This type of attack involves generating multiple paraphrases for each sentence, and choosing the one that increases the likelihood of disrupting statistical signatures embedded in the text, especially for token-level watermarking algorithms (Hou et al., 2023). While this type of attack significantly impacts most other models, SimMark demonstrates greater robustness against it. We must highlight that direct comparisons to $k$-SemStamp are not entirely fair, as $k$-SemStamp relies on domain-specific clustering of semantic spaces, making it domain-dependent. In contrast, SimMark, similar to SemStamp, is domain-independent. Still, SimMark outperforms $k$-SemStamp in most cases across various datasets and metrics, further underscoring its universality and robustness.
4.3 Paraphrase+Drop Attack
Figure 3: Detection results (ROC-AUC) under the Paraphrase+Drop Attack. Pegasus is used as the paraphraser. Here, 0 probability means no sentence is dropped, i.e., sentences are only paraphrased. SimMark is more robust than SemStamp and $k$-SemStamp across varying sentence drop probabilities.
To further evaluate the robustness of our method and its real-world applicability, we introduce a more challenging attack scenario, termed the Paraphrase+Drop Attack. We argue that an adversary not only paraphrases the text but also drops sentences deemed redundant. This reflects common editing practices since LLMs are known for their tendency to generate lengthy outputs with redundancy, likely due to their imperfect reward modelling (Chiang & Lee, 2024).
To simulate this scenario, we first paraphrase the text using the method described in the previous subsection and then randomly drop sentences from the paraphrased text. This combined attack tests the resilience of SimMark under more aggressive transformations that mimic real-world adversarial editing.

### v2

@@ 3.2 Watermarked Text Detection
The detection of SimMark follows a similar methodology to KGW by employing a $z$-test for hypothesis testing. However, akin to Hou et al. (2024a), the detection operates at the sentence level rather than the token level. To perform detection, we first divide the input text into sentences and use the same semantic embedding model to compute the embeddings for each sentence. If PCA was applied during the watermarking process, it must also be applied during detection to ensure consistency.
Algorithm 2 SimMark Detection Pseudo-Code
0: input text, embedding model, interval $[a,b]$, decay factor $K$, threshold $\beta$
1: Split the input text into sentences excluding the first sentence (i.e, the prompt): $M=\{X_{2},\dots,X_{N_{\mathrm{total}}}\}$, and set $N\leftarrow|M|$.
@@
8: Soft count of valid sentences: $N_{\mathrm{valid\_soft}}\leftarrow\sum_{i=2}^{N+1}c_{i}$.
@@
9: Estimate $p_{0}$ as the area under the human-written text embeddings similarity distribution curve within $[a,b]$.
10: Compute $z_{\mathrm{soft}}$ using Eq. (2).
11:if $z_{\mathrm{soft}}>\beta$ then
@@
12:  Reject $H_{0}$, i.e., text is likely generated by an LLM watermarked by SimMark.  [Uncaptioned image]
13:else
14:  Accept $H_{0}$, i.e., text is likely human-written.
15:end if
@@
Next, we compute the similarity ($s_{i+1}=sim(e_{i},e_{i+1})$) between consecutive sentences ($X_{i}$ and $X_{i+1}$) and count the number of valid sentences ($N_{\mathrm{valid\_soft}}$). Sentence $X_{i+1}$ is deemed valid if $s_{i+1}$ lies within the predefined interval $[a,b]$. However, paraphrasing may alter embeddings significantly, causing the similarity to deviate from the desired interval. To mitigate this, we adopt a soft counting approach, where a sentence is considered partially valid if its similarity is near the interval boundaries. Specifically, the soft count of $X_{i+1}$, denoted as $c_{i+1}$, is defined as follows:
$\begin{aligned} c_{i+1}=c(s_{i+1})=\begin{cases}1&\text{if }s_{i+1}\;{\color[rgb]{0,0.5,0}\in[a,b]},\\ e^{-K\min\{|a-s_{i+1}|,|b-s_{i+1}|\}}&{\color[rgb]{1,0,0}\text{otherwise.}}\end{cases}\end{aligned}$
@@
Here, $K>0$ is a decay factor controlling the smoothness of the soft counting. A higher $K$ makes the function behave closer to a step function, while a lower $K$ allows for smoother transitions, tolerating minor deviations outside the interval $[a,b]$. The total number of valid sentences is then computed as $N_{\mathrm{valid\_soft}}=\sum_{i}c_{i}$. Refer to Appendix I for an ablation study on how this approach improves robustness against paraphrasing, with only a minimal impact on performance in non-paraphrased scenarios, by allowing for some degree of error.
@@
During our initial experiments, we observed that, contrary to cosine similarity, Euclidean distance is very sensitive to paraphrasing, and even a subtle change in the sentences would result in a huge difference in the distances of embeddings. We hypothesize that Euclidean distance is sensitive to noise in high-dimensional spaces such as the semantic space of an embedder. To mitigate this, we propose using PCA: We fit a PCA model on a dataset of human-written texts to find the principal components of the sentence embeddings, i.e., the components that contribute the most to the semantical representation of the sentences. Then, we apply PCA to reduce embeddings’ dimension. More details on the dimensionality reduction are provided in Section 4 and Appendix J.
This approach can make reverse-engineering more difficult, as it would require knowledge of not only the embedder model, but also the PCA setting (e.g., number of components, access to the dataset used for fitting it, etc.). Without all these details, reproducing the similarity distribution becomes less straightforward.
The null hypothesis $H_{0}$ is defined as follows:
@@
$H_{0}$: The sentences are written by humans, i.e., the text sequence is generated without knowledge of the valid interval in the similarity of sentence embeddings.
@@
We calculate the $z$-statistic for the one-proportion $z$-test using the sample proportion $p=\frac{N_{\mathrm{valid\_soft}}}{N}$, where $N$ is the total number of samples (sentences). Since $N_{\mathrm{valid\_soft}}$ is a soft count of valid sentences, we refer to it as soft-$z$-score, which is given by $z_{\mathrm{soft}}=\frac{p-p_{0}}{\sqrt{\frac{p_{0}(1-p_{0})}{N}}}$ or alternatively:
@@
Here, the population proportion $p_{0}$ represents the ratio of valid sentences to all sentences in human-written text (i.e., a text with no watermark), which is estimated as the area under the similarity distribution curve of consecutive human-written sentences within the interval $[a,b]$ (like the one in Figure 9 in Appendix K). The value of $z_{\mathrm{soft}}$ can be interpreted as a normalized deviation of the number of valid sentences $N_{\mathrm{valid\_soft}}$ from its expectation $p_{0}N$.
As highlighted in Algorithm 2, the null hypothesis $H_{0}$ is rejected if $z_{soft}>\beta$, where $\beta$ is a threshold determined empirically by running the detection algorithm on human-written text. The threshold $\beta$ is selected to maintain a desired FP rate (i.e., minimizing the misclassification of human-written text as LLM-generated). Details on the computation of $\beta$ are provided in Appendix L.
4 Experiments & Results
@@
Table 1: Performance of different algorithms across datasets and paraphrasers, evaluated using ROC-AUC $\uparrow$ / TP@FP=1% $\uparrow$ / TP@FP=5% $\uparrow$, respectively (↑: higher is better). In each column, bold value indicates the best performance for a given dataset and metric, while underlined value denotes the second-best. SimMark consistently outperforms or is on par with other state-of-the-art methods across datasets, paraphrasers, and is the best on average.
Algorithm PPL $\downarrow$ Ent-3 $\uparrow$ Sem-Ent $\uparrow$
No watermark 11.89 11.43 3.32
UW 14.57 11.47 3.33
KGW 14.92 11.32 2.95
SIR 20.34 11.57 3.18
SemStamp 12.89 11.50 3.32
$k$-SemStamp 11.82 11.48 3.32
SimMark (ours) 12.69 11.50 3.37
Table 2: Comparison of the quality of text watermarked using different algorithms on the BookSum dataset (↓: lower is better, ↑: higher is better). SimMark yields quality metrics comparable to the no-watermark baseline, indicating minimal impact on text quality and semantic diversity. In contrast, token-level methods (UW, KGW, and SIR) notably degrade the text quality, especially in terms of perplexity.
Performance of SimMark is evaluated across different datasets and models using the area under the receiver operating characteristic curve (ROC-AUC) and true positive rate (TP) at fixed FP rates of 1% and 5% (TP@1%FP and TP@5%FP). Higher values indicate better performance across all metrics [^4] Like Hou et al. (2024a), all results are from a single run..
For dimensionality reduction, we fitted a PCA model on 8000 samples from the RealNews subset of the C4 dataset Raffel et al. (2020), reducing embedding dimensions from 768 to 16. After testing various principal component counts (ranging from 512 to 16), we found 16 to yield the best results. During our experiments, we evaluated both settings (with and without PCA). Specifically, PCA improved robustness against paraphrasing attacks when Euclidean distance was used (except on the BookSum dataset), but consistently degraded performance when cosine similarity was employed across all datasets. The results of these experiments are summarized in Table 7 in Appendix J.
Across all experiments, the decay factor was set to $K=250$, as this value provided an optimal trade-off between performance under both non-paraphrased and paraphrased conditions (see Appendix I for an ablation study on this). The threshold $\beta$ was determined empirically during the detection (refer to Appendix L for details) to achieve the specified FP rates (1% or 5%). The intervals $[0.68,0.76]$ for cosine similarity and $[0.28,0.36]$ for Euclidean distance with PCA and $[0.4,0.55]$ for Euclidean distance without PCA were found to be near-optimal, as detailed in Appendix K.
While the intervals as well as other hyperparameters such as decay factor $K$ could have been further optimized for each dataset or paraphraser individually, we chose not to do so to show the general performance of our method (in contrast, $k$-SemStamp fine-tunes domain-specific embedders that are optimized per setting).
@@ 4.1 Models and Datasets
For our experiments, we used the same fine-tuned version of OPT-1.3B Zhang et al. (2022) as in Hou et al. (2024a, b) [^5] Used AbeHou/opt-1.3b-semstamp (1.3B) model. to ensure fair comparison. However, we emphasize that our method is model-agnostic and it treats the LLM as a black-box text generator. As such, if the method performs well on one family of models, it is expected to generalize to others. To support this, we also tested our method on Gemma3-4B model Team et al. (2025) and observed similar results (see Appendix D). For semantic embedding, we utilized Instructor-Large [^6] Used hkunlp/instructor-large (335M) model. Su et al. (2023). Appendix B includes additional details on the experimental configurations.
In our experiments, we used three English datasets: RealNews subset of C4777Dataset card: allenai/c4 (Validation split) Raffel et al. (2020), BookSum [^8] Dataset card: kmfoda/booksum (Validation split) Kryscinski et al. (2022), and Reddit-TIFU [^9] Dataset card: ctr4si/reddit_tifu (Train split, short subset) Kim et al. (2019) datasets, as in Hou et al. (2024a). Specifically, 1000 samples from each dataset were chosen to analyze the detection performance and the text quality. Each sample was segmented into sentences [^10] Using sent_tokenize method of NLTK Bird et al. (2009)., with the first sentence serving as the prompt to the LLM.
@@
  • Tri-gram Entropy (Ent-3) ↑ Zhang et al. (2018): Assesses textual diversity via the entropy of the tri-grams distribution.
  • Semantic Entropy (Sem-Ent) ↑ Han et al. (2022): Measures semantic informativeness and diversity of the text.
  • Perplexity (PPL) ↓ Jelinek et al. (1977): Measures how surprising the text is to an oracle LLM [^11] Used facebook/opt-2.7b, following Hou et al. (2024a, b)..
4.2 Paraphrase Attack
@@
To evaluate the robustness of SimMark against paraphrase attacks, we tested it using three paraphrasers: $I.$ Pegasus paraphraser [^12] Used tuner007/pegasus_paraphrase (568M) model. Zhang et al. (2020), II.II. Parrot paraphraser [^13] Used parrot_paraphraser_on_T5 (220M) model. Damodaran (2021), $III.$ GPT-3.5-Turbo OpenAI (2022). Refer to Appendix N to find the prompts used with GPT-3.5-Turbo. Kirchenbauer et al. (2024) observed that prompting models to paraphrase entire texts often results in summarized outputs, with the summarization ratio worsening for longer inputs. To prevent any information loss caused by such summarization, we adopted a sentence-by-sentence paraphrasing scheme, which also ensures our results are comparable to Hou et al. (2024a, b). The quality of paraphrases was assessed using BertScore [^14] Used deberta-xlarge-mnli (750M) He et al. (2021). Zhang* et al. (2020), with all settings consistent with Hou et al. (2024a, b). The bottom part of Figure 2 demonstrates the paraphrase attack and detection phase in more detail.
@@
We also included the results for the bigram paraphrase attack introduced by Hou et al. (2024a), with identical settings (25 rephrases for each sentence when using Pegasus and Parrot, etc.). This attack involves generating multiple paraphrases for each sentence, and choosing the one that increases the likelihood of disrupting statistical signatures embedded in the text, especially for token-level algorithms Hou et al. (2024a). While this attack significantly impacts most other methods, SimMark demonstrates greater robustness against it. We must highlight that $k$-SemStamp relies on domain-specific clustering of semantic spaces, making it domain-dependent. In contrast, both SimMark and SemStamp are domain-independent. Still, SimMark outperforms both in nearly all cases across various datasets and metrics, further underscoring its universality and robustness.
4.3 Robustness to Sentence-Level Perturbations
To further evaluate the robustness of our method and its real-world applicability, we introduce more challenging attack scenarios: Paraphrase+Drop and Paraphrase+Merge Attacks. These scenarios simulate realistic adversarial editing strategies where a user attempts to remove or merge sentences after paraphrasing, while maintaining fluency.
Paraphrase+Drop assumes that an adversary not only paraphrases the text but also drops sentences deemed redundant. This reflects common editing practices, especially since LLMs often produce verbose outputs due to imperfect reward modeling Chiang and Lee (2024). To simulate this, we first paraphrase the input text and then randomly drop sentences with a specified probability $p$. Similarly, Paraphrase+Merge is designed to test robustness under more subtle structural changes. After paraphrasing, we replace end-of-sentence punctuations (e.g., ., ?, !) with the word “and” with probability $0<p<\frac{1}{2}$. In our experiments, we avoid higher values of $p$ as they result in unnaturally long and less fluent sentences. This setup simulates a realistic scenario where an adversary attempts to merge sentences while preserving the overall coherence of the text. These combined attacks serve to stress-test the resilience of watermarking methods under more naturalistic adversarial conditions that go beyond simple paraphrasing.